OpenStudy (anonymous):

Divide (2x^2+16x+21) ÷ (x+6) Your answer should give the quotient and the remainder.

OpenStudy (anonymous):

Look up division of polynomials :) quick youtube videos are easy!

OpenStudy (anonymous):

use... synthetic division

OpenStudy (anonymous):

i got the asnwer but not sure

OpenStudy (anonymous):

remainder is -3?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

whats the quotient

OpenStudy (anonymous):

The quotient is the solution more or less. Like 10+1=11. 11 is your sum

OpenStudy (anonymous):

2x+4?

OpenStudy (jdoe0001):

\(\large {\begin{array}{llccclllll} -6&|&2&+16&+21\\ &|&&-12&-24\\ \hline\\ &&2&4&3 \end{array}\implies 2x^2+16x+21 \div (x+6)\\ \quad \\ \color{blue}{2x+4+\frac{3}{x+6}}}\)

OpenStudy (jdoe0001):

for the quotient, you "drop" the original polynomial by "1 degree"

OpenStudy (anonymous):

so iwas right?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

hmmm actually... one sec

OpenStudy (jdoe0001):

+21 - 24 = -3 so....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea i got that remainder

OpenStudy (jdoe0001):

\(\large {\begin{array}{llccclllll} -6&|&2&+16&+21\\ &|&&-12&-24\\ \hline\\ &&2&4&-3 \end{array}\implies 2x^2+16x+21 \div (x+6)\\ \quad \\ \color{blue}{2x+4-\frac{3}{x+6}}}\)

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

quotient is 2x+4?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

the version I posted there is just the expression itself in full.... but yes, in terms of the division, if you keep the remainder, the quotient is just 2x +4

OpenStudy (anonymous):

TY

OpenStudy (jdoe0001):

yw